Determine the derivative of the function "f(x) = e"^("sqrt(x" ))


1
Expert's answer
2021-06-07T17:18:50-0400

"f(x)=e^{\sqrt{x}}\\
f'(x)=e^{\sqrt{x}}\frac{d}{dx}(\sqrt{x})\space(\text{by chain rule})\\
=e^{\sqrt{x}}\frac{1}{2\sqrt{x}}"